# Distributed optimization: designed for federated learning

### v2

@@ Abstract
Index Terms:
federated learning, data-driven, augmented Lagrangian, distributed optimization
@@ I Introduction
Another class of methods for solving problem (1) is based on the Alternating Direction Method of Multipliers (ADMM). The core idea is to introduce a consensus constraint into the primal problem and perform variable updates within the ADMM framework. Representative works in this direction include [28, 29, 30, 31, 32, 33, 34, 35, 36]. It is worth noting, however, that although these methods exhibit decentralized features in their problem formulation, their update mechanisms are essentially constrained by the classical two-block structure of ADMM. More precisely, they can be regarded as employing a two-level coordination scheme dominated by a central coordinator. Such a structure, as illustrated by the communication topology in Fig. 1(a), limits the flexibility and robustness expected in fully distributed systems.
FL, as an emerging privacy-preserving paradigm of distributed ML, faces critical challenges such as statistical heterogeneity and privacy concerns [37, 38]. Depending on communication topologies, FL can be broadly categorized into centralized and decentralized architectures [39], as illustrated in Fig. 1. Notably, existing FL algorithms predominantly adopt centralized paradigms, with Federated Averaging (FedAvg) emerging as the de facto standard due to its simplicity and low communication overhead [40]. Nevertheless, in scenarios with significant statistical heterogeneity—where data distributions vary considerably across clients—FedAvg is prone to “client drift” during training, resulting in unstable convergence and degraded performance [41]. To address this issue, FedProx was proposed as a representative alternative [42]. Although FedProx claims to mitigate the adverse effects of statistical heterogeneity, our theoretical analysis and empirical evaluations under various non-independent and identically distributed (non-IID) settings reveal that it remains inadequate in stabilizing convergence and preserving performance.
Motivated by the limitations of existing approaches—where DGD-based distributed optimization algorithms typically rely on assumptions of unbiased estimation [43, 44] or neglect such assumptions in their modeling, and consensus-based ADMM variants are structurally confined to two-level coordination—this study aims to develop a novel class of distributed optimization methods to address statistical heterogeneity while accommodating arbitrary communication topologies, thereby fulfilling the specific requirements of FL systems. The principal contributions are summarized as follows:
@@ II-A Related Work
where the symbol $\circ$ represents the Hadamard product in (4), i.e., $\mathbf{c}\circ\mathbf{d}=\left[c_{1},c_{2},\dots,c_{n}\right]^{\top}\circ\left[d_{1},d_{2},\dots,d_{n}\right]^{\top}=\left[c_{1}d_{1},c_{2}d_{2},\dots,c_{n}d_{n}\right]^{\top},$ and $\|\cdot\|$ denotes the 2-norm. To tackle the problem (2), we can leverage the standard augmented Lagrangian method.
@@ III Distributed Augmented Lagrangian Decomposition for Federated Learning
This section will introduces the Distributed Augmented Lagrangian Decomposition for FL (Fed-DALD) method, which is based on the coordination of subproblems across nodes and the construction of consensus constraints. These two aspects lead to the classification of the method into two types—centralized and decentralized—corresponding to the two communication topologies illustrated in Fig. 1.
@@ III-A Distributed Augmented Lagrangian Decomposition with Centralized Consensus
We formally introduce the first distributed optimization algorithm, termed Augmented Lagrangian Decomposition with Centralized Consensus for FL (Fed-DALD-CC). Its framework is outlined as follows.
@@ III-B Distributed Augmented Lagrangian Decomposition with Decentralized Consensus
Then, we introduce Lagrange multipliers $\mu=$ $\{\mu_{ij}\subseteq\mathbb{R}^{m}\mid i\in\mathbb{N}_{n},j\in\mathcal{R}_{i},j>i\}$ and the positive penalty $\rho$, consisting of elements from the set $\{\rho_{ij}\subseteq\mathbb{R}^{m}\mid i\in\mathbb{N}_{n},j\in\mathcal{R}_{i},j>i\}$.
@@ IV-A Convergence of Fed-DALD
    The Lagrangian (3) has a saddle point $(\mathbf{x}^{*},\mu^{*})\subseteq\mathbb{R}^{mn}\times\mathbb{R}^{q}$:
@@
Theorem 1 (Alternating Optimization) Assume (A1)(A2). Then, every limit point of $\{\mathbf{x}^{k,v}\}$ minimizes $\Lambda_{\rho}(\mathbf{x},\mu^{k})$ over $\mathbb{R}^{mn}$.
@@ V-A Topological Interpretation
In the hierarchical network, directed edges represent the flow of information, passing from lower-level clients to higher-level clients. A parent client can only start its computation after all its dependent lower-level clients have been completed and their information collected. The subproblem at the root node is the last one to be solved, while the nodes in each subsequent layer represent subproblems that need to be solved before those in higher layers. Fig. 2 shows some client coordination sequences represented by hierarchical networks, and Fig. 3 provides a detailed description of the corresponding sequences using hierarchical matrices.
@@ V-B2 Stochastic Optimization and Distributed Optimization for ML
$\mathcal{F}=\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})+\sum_{i=1}^{n}\nabla f_{i}(\mathbf{x}_{i})^{\top}(\hat{\mathbf{x}}-\mathbf{x}_{i})+\frac{1}{2\alpha}\sum_{i=1}^{n}\|\hat{\mathbf{x}}-\mathbf{x}_{i}\|^{2}.$ (45)
@@
$\left[\nabla_{i}^{2}f_{i}(\mathbf{x}_{i}^{k})-\frac{I}{\alpha}\right](\mathbf{x}_{i}^{k}-\hat{\mathbf{x}}^{k-1})=0.$ (46)
@@
$\hat{\mathbf{x}}^{k}=\mathbf{x}_{i}^{k}-\alpha\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (48)
@@
$\hat{\mathbf{x}}^{k}=\hat{\mathbf{x}}^{k-1}-\alpha\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (49)
@@
$\mathbf{x}_{i}^{k+1}=\mathbf{x}_{i}^{k}-\alpha\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (50)
Suppose we divide the $N$ samples into $n$ batches or clients, where each batch contains $N_{i}$ samples. Under the Fed-DALD-CC framework, implementing partial-cycle coordination, if we randomly select only one batch or client for update in the solving step (46), then algorithm (49) reduces to Mini-batch Gradient Descent (MBGD). Further, if we treat each batch or client as an independent system, and each sample as an individual client, with only one sample selected for optimization at each step (i.e., $B=1$), or when $N_{i}=1$, (49) further reduces to SGD. Similarly, by replacing the step size with $\alpha^{k}I=\nabla_{i}^{2}f(\mathbf{x}_{i}^{k})^{-1}$, we obtain the Stochastic Newton’s Method (SNM). Analogous to (41)–(44), it is straightforward to show that the above methods is convergent.
@@ V-B3 Inexact Version
$\mathbf{x}_{i}^{k+1}=\arg\min_{\mathbf{x}_{i}}\;f_{i}(\mathbf{x}_{i},\mathbf{x}^{k}_{j},\mathbf{x}^{k+1}_{e}).$ (57)
@@ VI-B Non-IID Case: Classification Training
Taking $\lambda=10^{-3}$ as an example, Table II demonstrates that, under the same number of iterations, the DALD algorithms significantly outperform FedProx. When the number of clients is $n=1000$ and the number of iterations reaches 3000, DALD-DC achieves an accuracy of 97.08%, representing an 11.77 percentage point improvement over FedProx’s 85.31%. As the number of clients increases from $n=10$ to $n=1000$, the accuracy of FedProx drops by 12.10% (from 97.41% to 85.31%), whereas DALD-DC exhibits only a 1.51% decrease (from 99.04% to 97.53%), indicating the superior adaptability of the proposed method in large-scale distributed scenarios.
Regarding stability, the standard deviation of FedProx increases from 28.23 at $n=10$ to 621.43 at $n=1000$, highlighting its limitations in handling statistical heterogeneity. In contrast, DALD-DC maintains a standard deviation of only 13.89 under the same conditions, demonstrating its effectiveness in suppressing performance fluctuations caused by statistical heterogeneity. Notably, as the number of clients grows, statistical heterogeneity becomes more pronounced. Table II reveals that for $n=500$ and $n=1000$, FedProx exhibits an order-of-magnitude increase in standard deviation, leading to significant performance degradation, whereas DALD-CC and DALD-DC maintain consistently low standard deviations with accuracy exceeding 95%, showcasing superior stability and generalization capability.
@@ VIII-B Computational Procedure for $\mathcal{D}^{\textit{k,v}}$ in Fed-DALD-DC
where $\{1\}$ (with a slight abuse of notation) represents the level identifier 1, corresponding to the first level in the hierarchical network. This also aligns with (70).

### v3

@@ I Introduction
Another class of methods for solving problem (1) is based on the Alternating Direction Method of Multipliers (ADMM). The core idea is to introduce a consensus constraint into the primal problem and perform variable updates within the ADMM framework. Representative works in this direction include [28, 29, 30, 31, 32, 33, 34, 35, 36]. It is important to note that, despite the decentralized characteristics exhibited by some of these methods in their formulations, all these works are, in essence, still limited by the classical two-block structure of ADMM. More precisely, they can be regarded as adopting a two-level coordination scheme governed by a central coordinator. As illustrated by the communication topology in Fig. 1(a), such a structure constrains the flexibility and robustness that are typically expected in fully distributed systems.
FL, as an emerging privacy-preserving paradigm of distributed ML, faces critical challenges such as statistical heterogeneity and privacy concerns [37, 38]. Depending on communication topologies, FL can be broadly categorized into centralized and decentralized architectures [39], as illustrated in Fig. 1. Notably, existing FL algorithms predominantly adopt centralized paradigms, with Federated Averaging (FedAvg) emerging as the de facto standard due to its simplicity and low communication overhead [40]. Nevertheless, in scenarios with significant statistical heterogeneity, where data distributions vary considerably across clients, FedAvg is prone to “client drift” during training, resulting in unstable convergence and degraded performance [41]. To address this issue, FedProx was proposed as a representative alternative [42]. Although FedProx claims to mitigate the adverse effects of statistical heterogeneity, our theoretical analysis and empirical evaluations under various non-independent and identically distributed (non-IID) settings reveal that it remains inadequate in stabilizing convergence and preserving performance.
Motivated by the limitations of existing approaches, where DGD-based distributed optimization algorithms typically rely on assumptions of unbiased estimation [43, 44] or neglect such assumptions in their modeling, and consensus-based ADMM variants are structurally confined to two-level coordination, this study aims to develop a novel class of distributed optimization methods to address statistical heterogeneity while accommodating arbitrary communication topologies, thereby fulfilling the specific requirements of FL systems.
The principal contributions are summarized as follows:
@@ II-A Related Work
where the symbol $\circ$ represents the Hadamard product in (4), i.e., $\mathbf{c}\circ\mathbf{d}=\left[c_{1},c_{2},\dots,c_{n}\right]^{\top}\circ\left[d_{1},d_{2},\dots,d_{n}\right]^{\top}=\left[c_{1}d_{1},c_{2}d_{2},\dots,c_{n}d_{n}\right]^{\top},$ and $\|\cdot\|$ denotes the 2-norm. To tackle the problem (2), we can leverage the standard augmented Lagrangian method presented in Algorithm 1.
@@ III Distributed Augmented Lagrangian Decomposition for Federated Learning
This section will introduces the Distributed Augmented Lagrangian Decomposition (DALD) method for FL, which is based on the coordination of subproblems across nodes and the construction of consensus constraints. These two aspects lead to the classification of the method into two types—centralized and decentralized—corresponding to the two communication topologies illustrated in Fig. 1.
@@ III-A Distributed Augmented Lagrangian Decomposition with Centralized Consensus
We formally introduce the first distributed optimization algorithm, termed Distributed Augmented Lagrangian Decomposition with Centralized Consensus for FL (Fed-DALD-CC). Its framework is outlined as follows.
@@ III-B Distributed Augmented Lagrangian Decomposition with Decentralized Consensus
Then, we introduce Lagrange multipliers $\mu=$ $\{\mu_{ij}\in\mathbb{R}^{m}\mid i\in\mathbb{N}_{n},j\in\mathcal{R}_{i},j>i\}$ and the positive penalty $\rho$, consisting of elements from the set $\{\rho_{ij}\in\mathbb{R}^{m}\mid i\in\mathbb{N}_{n},j\in\mathcal{R}_{i},j>i\}$.
@@ IV-A Convergence of Fed-DALD
    The Lagrangian (3) has a saddle point $(\mathbf{x}^{*},\mu^{*})\in\mathbb{R}^{mn}\times\mathbb{R}^{q}$:
@@
Theorem 1 Assume (A1)(A2). Then, every limit point of $\{\mathbf{x}^{k,v}\}$ minimizes $\Lambda_{\rho}(\mathbf{x},\mu^{k})$ over $\mathbb{R}^{mn}$.
@@ V-A Topological Interpretation
In the hierarchical network, directed edges represent the flow of information, passing from lower-level clients to higher-level clients. A parent client can only start its computation after all its dependent lower-level clients have been completed and their information collected. The subproblem at the root node is the last one to be solved, while the nodes in each subsequent level represent subproblems that need to be solved before those in higher levels. Hence, $j>i$ in (16) indicates that subproblem $j$ resides at a higher hierarchical level than subproblem $i$, and the two are interconnected within the hierarchical network. Fig. 2 shows some client coordination sequences represented by hierarchical networks, and Fig. 3 provides a detailed description of the corresponding sequences using hierarchical matrices.
@@ V-B2 Stochastic Optimization and Distributed Optimization for ML
$\mathcal{F}=\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})+\sum_{i=1}^{n}\nabla f_{i}(\mathbf{x}_{i})^{\top}(\hat{\mathbf{x}}-\mathbf{x}_{i})+\frac{1}{2}\sum_{i=1}^{n}\frac{\|\hat{\mathbf{x}}-\mathbf{x}_{i}\|^{2}}{\alpha_{i}}.$ (45)
@@
$\left[\nabla_{i}^{2}f_{i}(\mathbf{x}_{i}^{k})-\frac{I}{\alpha_{i}}\right](\mathbf{x}_{i}^{k}-\hat{\mathbf{x}}^{k-1})=0.$ (46)
@@
$\hat{\mathbf{x}}^{k}=\mathbf{x}_{i}^{k}-\alpha_{i}\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (48)
@@
$\hat{\mathbf{x}}^{k}=\hat{\mathbf{x}}^{k-1}-\alpha_{i}\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (49)
@@
$\mathbf{x}_{i}^{k+1}=\mathbf{x}_{i}^{k}-\alpha_{i}\nabla f_{i}(\mathbf{x}_{i}^{k}).$ (50)
Suppose we divide the $N$ samples into $n$ batches or clients, where each batch contains $N_{i}$ samples. Under the Fed-DALD-CC framework, implementing partial-cycle coordination, if we randomly select only one batch or client for update in the solving step (46), then algorithm (49) reduces to Mini-batch Gradient Descent (MBGD). Further, if we treat each batch or client as an independent system, and each sample as an individual client, with only one sample selected for optimization at each step (i.e., $B=1$), or when $N_{i}=1$, (49) further reduces to SGD. Similarly, by replacing the step size with $\alpha_{i}^{k}I=\nabla_{i}^{2}f(\mathbf{x}_{i}^{k})^{-1}$, we obtain the Stochastic Newton’s Method (SNM). Analogous to (41)–(44), it is straightforward to show that the above methods is convergent.
@@ V-B3 Inexact Version
$\mathbf{x}_{i}^{k+1}=\arg\min_{\mathbf{x}_{i}}\;f_{i}(\mathbf{x}_{i},\mathbf{x}^{k}_{j},\mathbf{x}^{k+1}_{e}),$ (57)
@@ VI-B Non-IID Case: Classification Training
Taking $\lambda=10^{-3}$ as an example, Table II demonstrates that, under the same number of iterations, the DALD algorithms significantly outperform FedProx. When the number of clients is $n=1000$ and the number of iterations reaches 3000, DALD-DC achieves an accuracy of 97.08%, representing an 9.87 percentage point improvement over FedProx’s 87.21%. As the number of clients increases from $n=10$ to $n=1000$, the accuracy of FedProx drops by 10.19% (from 97.40% to 87.21%), whereas DALD-DC exhibits only a 1.93% decrease (from 99.01% to 97.08%), indicating the superior adaptability of the proposed method in large-scale distributed scenarios.
Regarding stability, the standard deviation of FedProx increases from 32.12 at $n=10$ to 507.81 at $n=1000$, highlighting its limitations in handling statistical heterogeneity. In contrast, DALD-DC maintains a standard deviation of only 10.93 under the same conditions, demonstrating its effectiveness in suppressing performance fluctuations caused by statistical heterogeneity. Notably, as the number of clients grows, statistical heterogeneity becomes more pronounced. Table II reveals that for $n=500$ and $n=1000$, FedProx exhibits an order-of-magnitude increase in standard deviation, leading to significant performance degradation, whereas DALD-CC and DALD-DC maintain consistently low standard deviations with accuracy exceeding 95%, showcasing superior stability and generalization capability.
@@ VIII-B Computational Procedure for $\mathcal{D}^{\textit{k,v}}$ in Fed-DALD-DC
where $\{1\}$ (with a slight abuse of notation) represents the level identifier 1, corresponding to the bottom level in the hierarchical network. This also aligns with (70).